Ridgewood Company has 25,000 shares of cumulative 2% preferred stock, $50 par and 150,000 shares of $25 par common stock. During the current year, $62,500 was distributed as dividends. Determine the numerator in the dividend yield calculation.
Answer #1

Divided yield=Annual Divided per share÷Current Market Price

TOTAL DIVIDED = 62500

- Preference divided (25000*50*2%). =(25000)

Divided for equity holders = 37500

Divided per share = 37500÷150000 =. $0.25